# HiRT: Enhancing Robotic Control with Hierarchical Robot Transformers

### v1

@@ 2 Related Works
Vision-Language Models for Robotics. Applying pre-trained VLMs  [3, 4, 5, 6] to various embodied scenarios is a recent focal area of research. Most of the prior works focus on using VLMs for high-level planning or reasoning [27, 32, 33, 34, 35, 36, 37]. To effectively connect visual or linguistic information with the physical environment, embodied models need to fine-tune pre-trained VLMs on embodied data [1] including video data containing task-level planning in linguistic form [38, 17, 27], simple text descriptions [39, 40], low-level actions [41, 42, 43] (known as vision-language-action models). However, deploying such large VLA models often results in slow inference speeds [44], which makes embodied models unsuitable for scenarios requiring precise operations or quick execution. Our approach focuses on addressing this limitation by using a novel policy model, which can effectively retain the robust visual-linguistic capabilities of the larger models.
Hierarchical Action Planning. Hierarchical action planning [17, 45, 27, 46] involves decomposing a task into multiple simpler tasks that can be executed directly, enabling strategies to tackle more complex, long-horizon tasks. Previous works have demonstrated the role of inputting prompts into LLMs as a bridge to low-level actions. Specifically, this can be implemented through task-level planning [47, 38, 45], code execution [48, 49, 50], or other planning representations such as 3D scene graph [51], affordance function [52], and action pattern for locomotion [53]. However, these approaches are typically agnostic to physical embodiment, preventing the high-level models from directly interacting with the physical environment. In contrast to these methods, we ground VLMs to a specific robot’s physical form in an end-to-end manner, enabling it to learn hierarchical task planning through continuous intermediate representations.
@@ A.3 More Ablation
SafRT-AS 1 1.97Hz 96.7 77.2
SafRT 1 1.97Hz 94.2 76.0
SafRT-AS 6 13.42Hz 52.2 50.0
SafRT 6 13.42Hz 63.4 52.5
@@
To determine the impact of sampling images for VLM inputs, we compared the performance of SafRT with SafRT-AS under VLM settings of interval 1 and interval 6. As shown in Table 4, when VLM synchronizes with the action policy at every step during testing, SafRT-AS slightly outperforms SafRT. This is expected since SafRT-AS uses the most recently updated latent at each step during training. However, when VLM operates with an interval of 6 steps for inference, SafRT-AS shows nearly a 10% decrease in success rate compared to the original method, indicating that random sampling helps SafRT maintain better generalization performance during asynchronous operation.

### v3

@@ 2 Related Works
Vision-Language Models for Robotics. Applying pre-trained VLMs  [3, 4, 5, 6, 32] to various embodied scenarios is a recent focal area of research. Most of the prior works focus on using VLMs for high-level planning or reasoning [27, 33, 34, 35, 36, 37, 38]. To effectively connect visual or linguistic information with the physical environment, embodied models need to fine-tune pre-trained VLMs on embodied data [1] including video data containing task-level planning in linguistic form [39, 17, 27], simple text descriptions [40, 41], low-level actions [42, 43, 44] (known as vision-language-action models). However, deploying such large VLA models often results in slow inference speeds [45], which makes embodied models unsuitable for scenarios requiring precise operations or quick execution. Our approach focuses on addressing this limitation by using a novel policy model, which can effectively retain the robust visual-linguistic capabilities of the larger models.
Hierarchical Action Planning. Hierarchical action planning [17, 46, 27, 47, 48] involves decomposing a task into multiple simpler tasks that can be executed directly, enabling strategies to tackle more complex, long-horizon tasks. Previous works have demonstrated the role of inputting prompts into LLMs as a bridge to low-level actions. Specifically, this can be implemented through task-level planning [49, 39, 46], code execution [50, 51, 52], or other planning representations such as 3D scene graph [53], affordance function [54], and action pattern for locomotion [55]. However, these approaches are typically agnostic to physical embodiment, preventing the high-level models from directly interacting with the physical environment. In contrast to these methods, we ground VLMs to a specific robot’s physical form in an end-to-end manner, enabling it to learn hierarchical task planning through continuous intermediate representations.
@@ A.3 More Ablation
HiRT-AS 1 1.97Hz 96.7 77.2
HiRT 1 1.97Hz 94.2 76.0
HiRT-AS 6 13.42Hz 52.2 50.0
HiRT 6 13.42Hz 63.4 52.5
@@
To determine the impact of sampling images for VLM inputs, we compared the performance of HiRT with HiRT-AS under VLM settings of interval 1 and interval 6. As shown in Table 4, when VLM synchronizes with the action policy at every step during testing, HiRT-AS slightly outperforms HiRT. This is expected since HiRT-AS uses the most recently updated latent at each step during training. However, when VLM operates with an interval of 6 steps for inference, HiRT-AS shows nearly a 10% decrease in success rate compared to the original method, indicating that random sampling helps HiRT maintain better generalization performance during asynchronous operation.